Calculating Unemployment Given the following information about a country’s population, answer the questions which foll...

  1. Calculating Unemployment

Given the following information about a country’s population, answer the questions which follow. Make sure you show all of your calculations.

1987 1992.

Non-institutionalized, civilian,           182,753,000                191,576,000

Adult population

Employed persons                              112,440,000                117,598,000

Unemployed persons                          7,425,000                    9,384,000

  1. What was the employment rate in 1992? __________
  2. What was the labor force participation rate in 1987? ___________
  3. What was the labor force participation rate in 1992? ___________
  4. Explain how the unemployment rate can increase even if the number of people employed is increasing from 1987 to 1992?
Answer #1
  • Labor Force = Employed + Unemployed
  • LFPR = Labor Force / Civilian Non-Institutionalized Population
  • The unemployment rate is the percentage of unemployed workers in the labor force.1987 1992 Non-institionlized civilian 182,753,000 191,576,000 Employed 112, 440,000 117,598,000 Unemployed 7,425,000 9,384,00• In 1987: Labor force = = 112, 440,000 + 7, 425,000 119,865,000 000 -. LFPR = /119,865,000 182,753,000 = = 0.656 165.69 hen• In [1992] Labor force = 117,598,000 + 9,38 4,000 = 12 6,982,000 : LFPR = 126, 982,000 191,576,000 - 9384,000 URE (9389.)*10. Er in 1992 = 92.67. LFPR in 1987 = 65.67. LFPR in 1992= 66.37. unemployment rate increased from [6.2% in 1987 to 17.4% in 1